### v1

# SelfDefend: LLMs Can Defend Themselves against Jailbreaking in a Practical Manner

Abstract
Jailbreaking is an emerging adversarial attack that bypasses the safety alignment deployed in off-the-shelf large language models (LLMs) and has evolved into four major categories: optimization-based attacks such as the Greedy Coordinate Gradient (GCG) attack, jailbreak template-based attacks such as “Do-Anything-Now” (DAN), advanced indirect attacks like DrAttack, and multilingual jailbreaks. However, delivering a practical jailbreak defense is challenging because it needs to not only handle all the above jailbreak attacks but also incur negligible delay to user prompts, as well as be compatible with both open-source and closed-source LLMs.
Inspired by how the traditional security concept of shadow stacks defends against memory overflow attacks, this paper introduces a generic LLM jailbreak defense framework called SelfDefend, which establishes a shadow LLM defense instance to concurrently protect the target LLM instance in the normal stack and collaborate with it for checkpoint-based access control. The effectiveness of SelfDefend builds upon our observation that existing LLMs (both target and defense LLMs) have the capability to identify harmful prompts or intentions in user queries, which we empirically validate using the commonly used GPT-3.5/4 models across all major jailbreak attacks. Our measurements show that SelfDefend enables GPT-3.5 to suppress the attack success rate (ASR) by 8.97-95.74% (average: 60%) and GPT-4 by even 36.36-100% (average: 83%), while incurring negligible effects on normal queries. To further improve the defense’s robustness and minimize costs, we employ a data distillation approach to tune dedicated open-source defense models. These models outperform four SOTA defenses and match the performance of GPT-4-based SelfDefend, with significantly lower extra delays. We also empirically show that the tuned models are robust to targeted GCG and prompt injection attacks.
1 Introduction
Recent years have witnessed the significant potential of large language models (LLMs) in various domains [2], such as natural language processing (NLP) [3, 4, 5], information retrieval [6], image generation [7], science [8, 9, 10, 11], code tasks [12, 13, 14, 15], security tasks [16, 17, 18, 19, 20, 21], and more. To avoid causing social anxiety, ethical, and legal issues due to LLM responses to harmful questions, LLM vendors typically conduct safety alignment to prevent the misuse of LLMs through techniques like RLHF (Reinforcement Learning from Human Feedback) [22]. In response to a harmful prompt that violates safety policies, an aligned LLM often replies with a standard response such as “I’m sorry, I can’t assist with that request.” To bypass LLMs’ safety alignment checks, an adversarial attack known as jailbreaking [23] was proposed.
In the past two years, research on LLM jailbreak attacks and defenses has attracted considerable interest, with most of them focused on the offensive side. Specifically, jailbreak strategies have evolved from manual prompt engineering [23, 24, 25] to automatic LLM-based red teaming [26, 27, 28]. Besides these template-based jailbreaks aimed at identifying a valid jailbreak prompt template, a more generic, optimization-based adversarial jailbreak approach, notably Greedy Coordinate Gradient (GCG) [29], was recently proposed. It uses public available models to train adversarial suffixes that maximize the probability for LLMs to produce an affirmative response rather than refusing to answer. These studies [29, 30, 31] found that the identified suffixes are transferable to closed-source off-the-shelf LLMs. More recently, advanced indirect jailbreaks like DrAttack [32] and Puzzler [33], as well as multilingual jailbreaks [34, 35], have also been invented. In addition to proposing these four major types of jailbreaks, various benchmark studies on LLM jailbreak attacks [23, 36, 37, 38, 39, 40] have also been conducted.
On the contrary, the defensive side is somewhat overshadowed, despite around a dozen defense mechanisms being proposed in the past year. They can be roughly categorized into tuning-based and non-tuning-based mechanisms. Specifically, tuning-based defenses [41, 42, 43, 44, 45, 46] aim to fundamentally improve a model’s safety alignment against jailbreaking, while non-tuning-based defenses [47, 48, 49, 50, 51, 52, 53, 54, 55] can be typically applied to any off-the-shelf LLMs. We conduct an analysis of these defense techniques in §3 and find that it is still challenging for them to be widely used in practice. In short, we advocate that a practical jailbreak defense needs to not only handle all the aforementioned jailbreak attacks but also incur negligible delay to user prompts, as well as be compatible with both open-source and closed-source LLMs in various scenarios (see details in §3).
In this paper, we propose a new perspective on practically defending against jailbreak attacks, inspired by how the traditional security concept of shadow stacks [56] defends against memory overflow attacks. Similar to the shadow stack creating a shadow memory space, we establish a shadow LLM defense instance, $LLM_{defense}$, alongside the target LLM instance, $LLM_{target}$, in the normal stack. Under this framework, $LLM_{target}$ can process any user prompt query $P_{query}$ as usual to produce a token-by-token output. Meanwhile, $LLM_{defense}$ employs a dedicated detection prompt, $P_{direct}$ or $P_{intent}$, to wrap $P_{query}$ and detect its harmful prompt (via $P_{direct}$) or intention (via $P_{intent}$). Such a unique setup brings several benefits. ① It simultaneously utilizes both $LLM_{target}$’s own safety alignment and $LLM_{defense}$’s dedicated jailbreak detection, largely increasing the defense success rate due to this dual-layer protection. ② Since $LLM_{defense}$’s output is typically short, such as only “No” (indicating no issue) for normal queries, a checkpoint in the normal stack tends to be quickly triggered from the shadow stack without delaying $LLM_{target}$’s output. ③ Since $LLM_{defense}$ does not need to modify or monitor $LLM_{target}$’s internals, it can protect both open-source and closed-source LLMs. We concretize the above ideas into a generic jailbreak defense framework called SelfDefend and will introduce its details in §4.
The effectiveness of SelfDefend builds upon our observation that existing LLMs (both target and defense LLMs) have the capability to identify harmful portions (prompts/intentions) in user queries. To validate this hypothesis, we conduct an empirical measurement in §5 using the commonly used GPT-3.5/4 models under the SelfDefend architecture to test all major jailbreak attacks [29, 31, 36, 28, 32, 33, 34]. The results are quite promising in that SelfDefend enables both GPT-3.5 and GPT-4 to significantly suppress the attack success rate (ASR). Specifically, GPT-3.5-based SelfDefend reduces the ASR by 8.97-95.74% (average: 60%) compared to the baseline GPT-3.5, lowering the ASR to an average of 0.28, and GPT-4-based SelfDefend even reduces the ASR by 36.36-100% (average: 83%) compared to the baseline GPT-4, lowering the ASR to an extremely low average of 0.06. Besides the jailbreak scenarios, we also test SelfDefend against 805 normal prompts from the AlpacaEval dataset [57] and find that the pass rate is not affected for GPT-3.5 and slightly decreases by 3% for GPT-4, indicating that SelfDefend incurs negligible effects on normal queries. Moreover, SelfDefend incurs zero delay for over 95% of normal prompts across three out of four configurations, averaging 0.08 seconds for GPT-3.5 and 0.04 seconds for GPT-4. In various attack scenarios, SelfDefend causes an overall average extra delay of 0.04 seconds for GPT-3.5 and 0.39 seconds for GPT-4, respectively.
While the above measurements indicate that GPT-based SelfDefend effectively reduces the success rates of various types of jailbreak attacks, lowering the ASR to an average of 0.06, GPT-4 itself is commonly known to be expensive [58]. Moreover, the closed-source nature of GPT-3.5/4 raises privacy concerns for non-OpenAI model vendors. Therefore, we attempt to tune an open-source model that can be used under the SelfDefend framework for robust, low-cost, and self-contained jailbreak defense. By conducting GPT-4-based data distillation (with SelfDefend’s prompts) on a red-team dataset from Anthropic [59] comprising 38,961 harmful and harmless prompts, we generate a large set of high-quality tuning data. The distilled datasets are then used to tune our defense models through the LoRA fine-tuning technique [60] on the publicly available Llama-2-7b model [61]. The details will be introduced in §6.
To extensively evaluate our tuned models, we not only assess SelfDefend’s performance as in the aforementioned empirical measurement, but also compare with four representative jailbreak defenses: Perplexity Filter [62], SmoothLLM [51], Llama Guard [42], and Llama Guard 2 [63]. The results show that SelfDefend consistently outperforms all other defenses in 19 out of 21 tested attack scenarios (seven jailbreak methods and three target LLMs: GPT-3.5, GPT-4, and Llama 2) and reaches the level of GPT-4-based SelfDefend’s defense. The only two exceptions occurred in comparison with Perplexity Filter [62], which is specifically designed to defend against the GCG attack and offers no defense against other types of jailbreaks, particularly for the GCG attack targeting GPT-3.5 and Llama 2 LLMs. Nevertheless, we still outperform it when the target LLM is GPT-4.
Besides defense effectiveness, we also measure the extra delay $\Delta d$ caused by our defense models and find that the average $\Delta d$ is negligible, at 0-0.01 seconds for normal prompts. For attack scenarios, the maximum $\Delta d$ has decreased from 1.1 seconds in GPT-4-based SelfDefend to 0.25 seconds, with $\Delta d$ in all attack scenarios except for JSAA now below 0.1 seconds. These findings indicate that the tuning-based SelfDefend achieves negligible delays for both normal and jailbreak prompts, making it feasible for real-world deployment. Furthermore, we specifically assess whether the detected harmful portion actually aligns with the original prompt through the ensemble CLIP-score [64], and empirically show that the tuned models are robust to targeted GCG and prompt injection [65, 66] attacks that have white-box access to our defense models. More details are available in §7.
Availability and Ethics. To facilitate future research, we are on the way of releasing our datasets, raw evaluation results, and code at this anonymous GitHub link: https://github.com/SelfDefend. While our study required testing various jailbreak attacks, these tests never affected other OpenAI users and were conducted solely on our own OpenAI account and a local server. We also tried to minimize the overhead to OpenAI by spreading all the experiments over a period of two months instead of just several days and paid around 1,500 USD to OpenAI for our API usage.
2 Background
2.1 Threat Model
Attack Scenario. This research focuses on the attack scenario where an adversary seeks to perform jailbreaking on a text-based large language model (LLM). Multimodal jailbreaks [67, 68, 69, 70] are outside the scope of this paper. The objective of jailbreaking is to bypass the LLM’s safety mechanisms and induce it to generate harmful, toxic, or objectionable content. In this context, we assume the adversary has access to the LLM and can input arbitrary prompts to it. Given the vocabulary $\mathcal{T}$, the LLM, denoted as $\text{LLM}:\mathcal{T}^{\star}\to\Delta(\mathcal{T})$, takes a sequence of tokens $\mathcal{T}^{\star}$ as input and outputs a distribution $\Delta(\mathcal{T})$ over the next token in the sequence.
The adversary aims to find a prompt $P\in\mathcal{T}^{\star}$ that, when processed by the LLM, generates a response $R$ fulfilling a harmful goal $G$. We define a classifier $\text{JUDGE}:\mathcal{T}^{\star}\times\mathcal{T}^{\star}\to\{\text{True},\text{False}\}$, which returns True if and only if the response $R$ meets the criteria of the harmful goal $G$ given input $R,G$.
Adversary’s Objective. The adversary’s main objective is to generate responses from the LLM that are classified as successful jailbreaks. Specifically, the adversary seeks to maximize the probability that a response $R$ generated by the LLM for a given prompt $P$ is classified as harmful according to the goal $G$. Formally, the adversary’s objective can be expressed as:
$\sup_{P\in\mathcal{T}^{\star}}\Pr_{R\sim\text{LLM}(P)}[\text{JUDGE}(R,G)=\text{True}]$
where the randomness is due to the stochastic nature of the LLM’s responses to the input prompt $P$. In essence, the adversary iterates over potential prompts to find one that maximizes the likelihood of producing a harmful output as judged by the classifier.
Constraints and Assumptions. We assume the following constraints and assumptions in the general jailbreak threat model:
  • The adversary has black-box access to the LLM, meaning they can input prompts and observe outputs but do not have direct access to the model’s internal parameters or training data.
  • The goal string $G$ is predefined and represents a specific type of harmful content that the adversary aims to induce from the LLM.
  • The classifier JUDGE is accurate in determining whether a response constitutes a successful jailbreak according to the harmful goal $G$.
Under these assumptions, the adversary’s challenge is to discover prompts that effectively exploit the LLM’s vulnerabilities, leading to the generation of undesirable content while navigating the constraints of black-box access.
2.2 Jailbreak Attacks
Existing jailbreak attacks can be roughly grouped into four categories: optimization-based, template-based, indirect, and multilingual jailbreaks.
Optimization-based Jailbreak typically appends a suffix to a malicious prompt and iteratively updates this suffix using gradient-based or search-based methodologies for jailbreaking. The pioneering work by GCG [29] introduced a method known as the greedy coordinate gradient to optimize adversarial suffixes, facilitating transferable jailbreaks across various prompts and models. Sitawarin et al. [30] further extended this technique to GCG++ by employing a proxy model to guide the optimization process. Beyond gradient-based optimization, Andriushchenko et al. [31] utilized a simple random search on a suffix to maximize the target probability.
Template-based Jailbreak involves the use of prompt engineering [23, 24, 25] or LLM-based red teaming [26, 27, 36, 71, 28] to create effective prompt templates that can mislead LLMs into producing restricted content. Early template-based attacks combine manually crafted templates [23, 24, 25] with malicious instructions to achieve jailbreaking. For example, Liu et al. [24] proposed role-playing prompts to confuse LLMs. A more sophisticated approach, AutoDAN [28], employs genetic algorithms to optimize these hand-crafted templates. Recent advancements include the use of LLMs to assist in generating jailbreak prompts. PAIR [36] is the first method to utilize an auxiliary LLM to generate jailbreaks. Mehrotra et al. [71] then refined PAIR’s approach through a tree-based search method.
Indirect Jailbreak conceals malicious intents within the query text to circumvent the safety mechanisms of LLMs and elicit the desired malicious response. A straightforward method for executing indirect jailbreaks is by performing word substitution on the original malicious instruction [72]. Recently, DrAttack [32] introduced a technique that decomposes a malicious prompt into separate sub-prompts, effectively masking its underlying malicious intent. Puzzler [33] provides clues related to the malicious prompt, thereby guiding the target LLM towards a jailbreak.
Multilingual Jailbreak translates the harmful prompt into a language in which LLMs are less aligned for safety [34, 35]. Deng et al. [34] found that it is easier to jailbreak LLMs in low-resource languages than in high-resource languages and released a multilingual jailbreak prompt dataset, MultiJail. Besides the multilingual strategy, Wei et al. [23] encode the original harmful prompt into Base64 format to reduce the sensitivity of LLMs.
TABLE I: A comparison of existing jailbreak defenses under the four objectives we envisioned in §3.1. Note that the first six rows denote tuning-based defenses, while the last seven rows represent non-tuning-based defenses.
Venue1 Core Idea Objectives
O1 O2 O3 O4
Perplexity [41, 62] arXiv:2308.14132 Calculate the perplexity of the prompt to detect adversarial suffixes
Llama Guard [42] arXiv:2312.06674 Align the classification of both prompt and response
RPO [73], DPP [74] arXiv:2401.17263 Optimize universal and transferable suffixes that enforce safe outputs
DRO [43], PAT [44] SeT @ ICLR’24 Add a safety prompt and optimize it as a prefix using prompt tuning
SafeDecoding [45] arXiv:2402.08983 Compare and amplify the token probabilities of safety disclaimers
LED [46] arXiv:2405.18166 Identify safety-critical layers and realign them through model editing
Self Defense [47] arXiv:2308.07308 Add one more step to check the safety of original LLM responses
RAIN [48] ICLR’24 Self-evaluate each token output, rewind, and determine the final output
erase-and-check [49] arXiv:2309.02705 Erase some tokens from the prompt; check the rest using a safety filter
[50, 51, 52] arXiv:2309.14348 Perturb copies of each prompt and aggregate their output responses
IAPrompt [53] arXiv:2401.06561 Conduct a chain-of-thought intention analysis for each input prompt
GradSafe [54] arXiv:2402.13494 Compare the prompt’s gradient similarity with safety-critical gradients
Gradient Cuff [55] arXiv:2403.00867 Compare the gradient norm of refusal loss with that in benign queries
  • = applies; = partially applies; = does not apply.
  • 1
    We only present the venue information for the earliest article in each line of defense techniques.
3 Objectives and Related Work
3.1 Design Objectives
With various jailbreak attacks presented in §2, we now envision the design objectives of an ideal defense as follows:
O1
Handling all kinds of jailbreak attacks. The proposed defense should be able to handle all categories of jailbreak attacks listed in §2, including not only optimization-based adversarial jailbreaks but also template-based, indirect, and multilingual jailbreaks.
O2
Incurring negligible delay to user prompts. The proposed defense should not impact the user experience, causing either no delay or only a negligible one to normal user prompts.
O3
Providing explanations for potential jailbreak queries. Similar to O2, when the proposed defense detects any query potentially related to jailbreaking, it should provide helpful explanations on why the query is considered harmful.
O4
Compatible with both open-source and closed-source LLMs. The proposed jailbreak defense approach should protect both white-box, open-source LLMs as well as black-box, closed-source LLMs.
We clarify that compared with the other three objectives, O3 is not a mandatory objective. Nevertheless, we believe O3 is particularly important for users to forensically understand potential jailbreak queries and better protect against jailbreak attacks in the future.
3.2 Analysis of Existing Defenses
Table I summarizes our analysis of major LLM jailbreak defenses under the four objectives we envisioned above. They can be roughly categorized into tuning-based (the first six rows) and non-tuning-based (the last seven rows) mechanisms. Specifically, tuning-based defenses aim to fundamentally improve a model’s safety alignment against jailbreaking, while non-tuning-based defenses can be typically applied to any off-the-shelf LLMs to enhance their safety against jailbreak attacks. In the rest of this subsection, we present our analysis results from four perspectives:
First, most jailbreak defenses target multiple kinds of jailbreak attacks but typically do not cover advanced indirect attacks (O1: ), while a few existing defense mechanisms are specifically designed to defend against optimization-based adversarial attacks only (O1: ). The latter includes two perplexity-based filtering approaches [41, 62] and several input perturbation-based approaches [49, 50, 51, 52]. Specifically, Alon and Kamfonas [41] proposed the first tuning-based defense in the sense that they not only leveraged perplexity values as an indicator to detect prompts with adversarial suffixes but also tuned a classifier to consider both the sequence length of prompts and their perplexity for improved filtering. Another line of GCG-specific jailbreak defenses perturbed copies of the input prompt and aggregated the output responses, with SmoothLLM [51] as a representative example.
Second, the majority of prior defenses inherently incur additional delays to user prompts (O2: ), with only two exceptions. Since the design principles of most prior defenses are to check the target LLM’s responses [47, 48, 42] and internal statuses [41, 62, 45, 54, 55, 46], or to conduct extra-round analyses of the input prompt [49, 50, 51, 52, 53], it is thus inherently difficult for these approaches to avoid additional delay. The only exceptions are [73, 74, 43, 44], which conducted optimization-based suffix or prefix training for defensive purposes, with the training objective being the exact opposite of the GCG attack. As a result, the tuned models behave like normal LLMs, incurring no extra delay to user prompts.
Third, most non-tuning-based defenses have the potential to provide explanations for jailbreak queries (O1: or ), whereas half of the tuning-based defenses cannot because they rely solely on LLMs’ internal indicators for tuning (O3: ). To provide explanations for potential jailbreak queries, a defense mechanism needs to understand the semantics of incoming prompt queries. Hence, it is hard for approaches that rely solely on LLMs’ internal indicators, whether they are tuning-based [45, 46] or non-tuning-based [54, 55], to provide straightforward explanations to users.
Fourth, most non-tuning-based defenses are compatible with both open-source and closed-source LLMs (O4: ), except for those that need to leverage LLMs’ internal indicators, while the opposite is true for tuning-based defenses (O4: ). Unless they need to monitor LLMs’ internal indicators [48, 54, 55], non-tuning-based defenses, such as Self Defense [47] and IAPrompt [53], are naturally compatible with both open-source and closed-source LLMs. On the other hand, tuning-based defenses typically require white-box access to the LLMs to enable their in-depth defense. The only exception is Meta’s Llama Guard [42], which aligns the classifications of both prompt and response, and thus, theoretically, it can serve as a plug-in model to work with both open-source and closed-source LLMs.
Figure 1: A high-level overview of the SelfDefend framework and our study.
4 The SelfDefend Framework
After analyzing the pitfalls of existing defenses, we propose a new perspective on defending LLMs against jailbreaks. Our key idea is to deploy a dedicated LLM alongside the target LLM to concurrently detect potential jailbreak queries. This idea is made possible because we found that LLMs do have the capability to protect themselves by identifying harmful portions in user queries.
Based on this idea, we propose a generic LLM jailbreak defense framework called SelfDefend. As shown in Figure 1, SelfDefend creatively establishes a shadow stack alongside the normal stack in the LLM space to conduct checkpoint-based access control, which mimics traditional security defense concepts such as the shadow stack for defending against buffer overflow attacks [56] and the library-based checkpoint from SCLib [75]. We denote the target LLM in the normal stack as $LLM_{target}$ and the defense LLM in the shadow stack as $LLM_{defense}$. SelfDefend simultaneously utilizes both $LLM_{target}$’s own safety alignment and $LLM_{defense}$’s dedicated jailbreak detection, largely increasing the defense success rate. In SelfDefend, $LLM_{defense}$ can be instantiated from the same model as $LLM_{target}$, although in practice we suggest using a dedicated $LLM_{defense}$ that is robust and low-cost for detecting jailbreak queries. Since $LLM_{defense}$ does not need to modify $LLM_{target}$ or know any of its internal details, SelfDefend can protect all kinds of target LLMs, whether they are open-source or closed-source (O4: ).
Given an incoming prompt query $P_{query}$, SelfDefend dispatches it to both $LLM_{target}$ and $LLM_{defense}$ for concurrent processing. Specifically, $LLM_{target}$ processes $P_{query}$ as usual, whether it is a normal user prompt or an adversarial jailbreak prompt, but caches its token-by-token output until a checkpoint is triggered from the shadow stack. In contrast, $LLM_{defense}$ employs a dedicated detection prompt, $P_{direct}$ or $P_{intent}$, to wrap $P_{query}$ and detect its harmful prompt (via $P_{direct}$) or intention (via $P_{intent}$). Once a token of “No” (indicating no issue) or a harmful portion (prompt/intention) is output, $LLM_{defense}$ triggers a checkpoint in the normal stack so that $LLM_{target}$ can decide whether to release the token-by-token output. Since the time for outputting the token “No” is very limited, the additional delay introduced by SelfDefend is negligible for normal user prompts (O2: ). Moreover, the identified harmful portion can help generate an explainable LLM response to adversarial prompts (O3: ). These unique advantages make SelfDefend the first practical jailbreak defense compared to the existing defense mechanisms explained in §3.2.
Now the effectiveness of SelfDefend depends on whether it can handle all kinds of jailbreak attacks (i.e., O1). For this, we proceed to conduct an empirical study in §5 to validate our earlier hypothesis that “LLMs have the capability to identify harmful portions in user queries.”
5 An Empirical Measurement
In this section, we conduct extensive measurements to empirically show that existing LLMs, notably the commonly used GPT-3.5/4 models, perform acceptably under the SelfDefend framework. This not only validates the hypothesis assumed in §4 but also enables widespread, convenient usage of SelfDefend without the need for further tuning of a defense model. For example, various custom GPTs applications [76] can adopt SelfDefend to immediately enhance their defense capabilities against jailbreaks.
Figure 2: SelfDefend’s two detection prompt templates.
5.1 Initializing SelfDefend for GPT-3.5/4
To facilitate measurements, we need to first initialize SelfDefend for GPT-3.5/4. For GPT-3.5-based experiments, both $LLM_{target}$ in the normal stack and $LLM_{defense}$ in the shadow stack are set to GPT-3.5. Likewise, for GPT-4-based experiments, both $LLM_{target}$ and $LLM_{defense}$ are set to GPT-4.
Prompt Templates. $LLM_{target}$ directly uses user prompts $P_{query}$ but we need to set up the detection prompts for $LLM_{defense}$. As mentioned earlier in §4, we have designed $P_{direct}$ and $P_{intent}$ for two different defense schemes. Figure 2 shows their detailed prompt templates. We can see that $P_{direct}$ aims to directly pinpoint the harmful part from the original jailbreak/normal $P_{query}$; otherwise, it outputs a token of “No.” In contrast, $P_{intent}$ employs a chain-of-thought (CoT) [77] approach to first summarize the true intention of $P_{query}$ and then, based on the summarized intention, further extract any safety-violation part, if present. It also includes two examples in the prompts to help LLMs better understand the instructions and conduct few-shot in-context learning [78].
By comparing these two prompts, we expect that $P_{direct}$ would perform better on optimization-based jailbreaks that directly include the jailbreak portion, while $P_{intent}$ might be more effective against indirect jailbreaks. One additional objective of the empirical measurement in this section is to explore the best setup for deploying SelfDefend with off-the-shelf LLMs under four different combinations (GPT-3.5 vs. GPT-4; $P_{direct}$ vs. $P_{intent}$).
Environment Setup. In our implementation, we use the OpenAI API in Python to access GPT-3.5 and GPT-4. The API versions of GPT-3.5 and GPT-4 are gpt-3.5-turbo-0125 and gpt-4-0125-preview, respectively. We set the hyper-parameters max_tokens, temperature, n, and top_p to 128, 0, 1, and 0, respectively. When accessing the GPT APIs, no system prompt is added.
TABLE II: The details of our collected benchmark datasets.
Dataset # Prompts Jailbreak Methods
JailbreakBench [79] 100 GCG [29], JSAA [31]
PAIR [36], AutoDAN [28]
DrAttack [32], Puzzler [33]
MultiJail [34] 315 MultiJail
AlpacaEval [57] 805 Normal Prompts
TABLE III: The ASR results from testing LLMs against four major categories of jailbreak attacks and normal prompts.
LLMs Optimization-based Template-based Indirect Multilingual Normal
GCG JSAA PAIR AutoDAN DrAttack Puzzler MultiJail AlpacaEval
GPT-3.5 (baseline) 0.56 0.94 0.72 0.90 0.78 0.98 0.39 0.97
GPT-3.5-based Shadow Stack ($P_{direct}$) 0.72 0.98 0.77 0.96 1.00 0.98 0.87 0.97
GPT-3.5-based SelfDefend ($P_{direct}$) 0.45 0.92 0.60 0.87 0.78 0.96 0.36 0.95
GPT-3.5-based Shadow Stack ($P_{intent}$) 0.28 0.06 0.37 0.35 0.86 0.22 0.52 0.99
GPT-3.5-based SelfDefend ($P_{intent}$) 0.19 0.04 0.29 0.31 0.71 0.22 0.20 0.97
Reduction factor 66.07% 95.74% 59.72% 65.55% 8.97% 77.55% 48.71% 0%
GPT-4 (baseline) 0.08 0.16 0.33 0.19 0.74 0.90 0.08 0.97
GPT-4-based Shadow Stack ($P_{direct}$) 0.01 0.00 0.11 0.01 0.05 0.27 0.14 0.96
GPT-4-based SelfDefend ($P_{direct}$) 0.00 0.00 0.10 0.01 0.04 0.26 0.01 0.94
Reduction factor 100% 100% 36.36% 94.73% 94.59% 71.11% 87.50% 3.09%
GPT-4-based Shadow Stack ($P_{intent}$) 0.08 0.05 0.21 0.07 0.13 0.36 0.30 0.99
GPT-4-based SelfDefend ($P_{intent}$) 0.05 0.05 0.18 0.07 0.13 0.28 0.01 0.97
(a) The CDF plot of $\Delta d$ for normal prompts.
(b) Average $\Delta d$ for various jailbreak prompts.
Figure 3: The extra delay $\Delta d$ caused by SelfDefend for normal prompts and jailbreak prompts, respectively.
5.2 Datasets and Attack Setup
After SelfDefend is set up, we collect benchmark datasets, implement the corresponding attack methods, and determine the effectiveness metrics.
Benchmarks. Based on the four categories of existing jailbreak attacks we surveyed in §2 — optimization-based, template-based, indirect, and multilingual jailbreaks — we identify representative jailbreak attack methods in each category. We then collect three benchmark datasets, JailbreakBench [79], MultiJail [34], and AlpacaEval [57], from which we use their user prompts for testing SelfDefend. Table II lists the details of our collected benchmark datasets. Specifically, we use a set of 100 instructions from JailbreakBench [79], a standardized evaluation framework, to drive optimization-based jailbreaks (GCG [29] and JSAA [31]), template-based jailbreaks (PAIR [36] and AutoDAN [28]), and indirect jailbreak attacks (DrAttack [32] and Puzzler [33]). In contrast, we directly use the original prompts from the MultiJail and AlpacaEval datasets to construct inputs for the scenarios of multilingual jailbreaks and normal prompts, respectively.
Attack Setup. For GCG, we choose its individual version and optimize the suffix on Vicuna-7b-v1.3 [80] with a batch size of 512 and 500 optimization steps. For JSAA, we query GPT-3.5 with 10,000 iterations and 10 restarts. We perform transfer attacks on GPT-4 using the prompts generated on GPT-3.5. For PAIR, we select Vicuna-13b-v1.5 [80] as the attack model. For AutoDAN, we choose its first version of the genetic algorithm, i.e., AutoDAN-GA. The genetic algorithm used in AutoDAN-GA features a crossover rate of 0.5, a mutation rate of 0.01, and 100 optimization steps. For DrAttack and Puzzler, we use GPT-4 to construct their jailbreak prompts. For MultiJail, we selected all 315 queries in the Bengali language. For AlpacaEval, we select all 805 questions from the AlpacaEval dataset.
Metrics. We measure the defense effectiveness of SelfDefend indirectly using the attack success rate (ASR) [29] or unsafe rate [34]. Following [29, 28], ASR is measured by detecting whether the LLM response contains keywords indicating a refusal to answer, like “I can’t assist.” If such a keyword or its variant is included, it indicates that the attack is unsuccessful; otherwise, it indicates that the jailbreak is successful. We use the common practice [29] for a list of such keywords, which is available in Appendix A. GCG, JSAA, PAIR, AutoDAN, DrAttack, and Puzzler all adopt ASR as the metric for measuring jailbreak performance. The lower the ASR, the stronger the defense performance of SelfDefend against jailbreaks. Note that we also use ASR to evaluate the performance of SelfDefend under normal prompts from AlpacaEval, but a higher ASR indicates that our framework is more compatible with normal prompts (in such a case, ASR can be interpreted as answer success rate). On the other hand, the unsafe rate uses GPT-4 to determine whether the response of the target model matches the jailbreak goal, following [34]. We adopt the unsafe rate in a way similar to ASK to specifically measure the jailbreak performance of MultiJail, since the response to a multilingual prompt does not contain English keywords.
5.3 Measurement Results
Table III summarizes our core measurement results, which we separate into the defense effectiveness against jailbreaks and the impact on normal user prompts. Additionally, we measure the extra delay introduced by SelfDefend.
Baseline Performance. We first measure the baseline performance of using GPT-3.5/4 only under the tested jailbreaks. As shown in Table III, these jailbreaks are much more effective on GPT-3.5 compared to GPT-4. The ASR for GPT-3.5 can be as high as 0.39 to 0.98 (average: 0.75), whereas the ASR for GPT-4 typically ranges from 0.08 to 0.33, except for indirect jailbreaks in which case GPT-4 also has a high ASR (0.74 for DrAttack and 0.9 for Puzzler).
Defense Effectiveness. We then compare ASR after equipping SelfDefend with the baseline model. In ablation, we also report ASR using only the shadow stack. Our findings are summarized as follows:
  • Based on GPT-3.5, although the SelfDefend frameworks supported by intent prompt and direct prompt both decrease success rates, there is a considerable difference in the enhancement of defense performance achieved by the two methods. The SelfDefend framework with direct prompt shows a modest reduction in terms of ASR due to its internal shadow stack performing even worse than GPT-3.5. For example, its ASR on AutoDAN is 96%, compared to GPT-3.5’s 90%. In contrast, the SelfDefend framework utilizing intent prompt demonstrates significantly better performance than the one with direct prompt, with the reduction factor ranging from as high as 95.74% to as low as 8.97%.
  • Based on GPT-4, both SelfDefend frameworks significantly reduce the ASRs of various jailbreak attacks. Different from the results on GPT-3.5, SelfDefend with direct prompt is better than SelfDefend with intent prompt due to the difference between the two shadow stacks. Although jailbreak attacks present lower ASRs than on GPT-3.5, the reduction factor of SelfDefend with direct prompt ranges from 36.36% to 100.00%.
  • SelfDefend with direct prompt and SelfDefend with intent prompt present the best defense enhance on GPT-4 and GPT-3.5, respectively. Furthermore, SelfDefend with direct prompt demonstrates the higher reduction factor than SelfDefend with intent prompt on GPT-3.5. Thus, we contend that SelfDefend with direct prompt relies more on the model’s own discrimination ability so that is suitable for powerful models such as GPT-4, while the intent prompt can strengthen the model’s logical reasoning and still has decent defense capabilities even on weaker models.
Impact on Normal Queries. Besides defense effectiveness, we also measure the impact of SelfDefend on normal queries by evaluating the ASR of 805 normal prompts from AlpacaEval. Table III shows that for both GPT-3.5 and GPT-4, SelfDefend (with $P_{direct}$) slightly reduces the ASR by 2% (from 0.97 to 0.95) and 3% (from 0.97 to 0.94), respectively. Meanwhile, SelfDefend (with $P_{direct}$) does not decrease the ASR at all for either model. Considering that the ASR under SelfDefend is still very high and close to the baseline, we conclude that SelfDefend’s defense would not notably impact normal user queries.
Extra Delay $\Delta d$. Recall that one of our design objectives is to incur negligible delay to user prompts (O2 in §3.1). We thus further measure the extra delay $\Delta d$ caused by SelfDefend, which is calculated as follows:
[…]
where $d_{total}$ denotes the total delay to SelfDefend-protected LLMs, and $d_{normal}$ represents the separate time cost of target LLMs in the normal stack. The metric of $\Delta d$ is mainly important for normal prompts because it would directly affect the experience of normal users. Figure 3(a) plots the cumulative distribution function (CDF) of $\Delta d$ for normal prompts. We can see that over 95% of the cases incur zero delay across three configurations of SelfDefend, demonstrating that SelfDefend is capable of defending against jailbreaks at the cost of negligible delay for normal user prompts. The only exception occurs with SelfDefend using $P_{intent}$ and GPT-3.5, where around 80% of the cases have no extra delay.
For jailbreak prompts, we also calculate their average $\Delta d$ for different attack scenarios in Figure 3(b). We can see that GPT-3.5’s average $\Delta d$ is under 0.1 seconds for all eight scenarios, while GPT-4 incurs negligible $\Delta d$ only in the normal scenario of AlpacaEval. This is likely because the baseline GPT-4 has good defense performance for most jailbreaks, suggesting its original output is already short and the delay is minimal.
Figure 4: The training procedure for fine-tuning our defense model.
6 Tuning a Dedicated Defense Model
Based on the empirical measurement results, we have observed that GPT-4, when equipped with SelfDefend, significantly reduces the success rates of various types of jailbreak attacks, lowering the ASR to an average of 0.06 (for GPT-4-based SelfDefend with $P_{direct}$). However, GPT-4 is commonly known to be expensive [58], and its closed-source nature raises privacy concerns for non-OpenAI model vendors. Therefore, our objective in this section is to tune an open-source model that can be used under the SelfDefend architecture for robust, low-cost, and self-contained jailbreak defense.
6.1 Design Overview
Given the powerful defensive capability of GPT-4-based SelfDefend, our intuition is to “transfer” this capability to an open-source model. To do so, we leverage GPT-4-based SelfDefend to distill and generate high-quality tuning data. Figure 4 depicts our proposed pipeline to fine-tune an open-source defense model. Specifically, by continuously incorporating harmful and harmless prompts into our defense prompts (i.e., $P_{direct}$ or $P_{intent}$) as inputs for GPT-4, we gather their outputs as labels for these samples. Since we utilize two defense prompts, we eventually obtain two separate datasets, which we then use to fine-tune the employed open-source model for our defense purposes.
6.2 Data Distillation
To distill GPT-4’s “knowledge” on safety policies, a dataset containing both harmful and harmless queries is necessary. We utilize the red-team data from Anthropic [59] as the query dataset. This dataset consists of 38,961 text transcripts documenting conversations between human adversaries and AI assistants. From this dataset, we select the initial human prompt and exclude the corresponding assistant response, omitting any subsequent exchanges, to create a single-turn prompt dataset designated as $\mathcal{D}_{red}$. It is worth noting that this dataset is released in 2022 and has no overlap with JailbreakBench/MultiJail released in 2024. Furthermore, this dataset is exclusively in English, implying that a model trained on this data will not acquire any additional multilingual capabilities.
Next, we employ GPT-4 and our defense prompts to distill GPT-4’s “knowledge” on individual queries from $\mathcal{D}_{red}$. This procedure can be formalized in Equation 2. Note that, for simplicity, we refer to the symbols $P_{direct}$ and $P_{intent}$ as $P_{dir}$ and $P_{int}$, respectively. Their corresponding datasets are denoted as $D_{dir}$ and $D_{int}$, respectively.
$\mathcal{D}_{dir}$ =$=\{(x,y)|x\in\mathcal{D}_{red}\},\quad y=\operatorname{GPT_{4}}(P_{dir}[x])$ (2)
$\mathcal{D}_{int}$ =$=\{(x,y)|x\in\mathcal{D}_{red}\},\quad y=\operatorname{GPT_{4}}(P_{int}[x]),$
Here, $x$ denotes an original prompt from $\mathcal{D}_{red}$ and $y$ represents the corresponding response from GPT-4 under either $P_{dir}$ or $P_{int}$. By analyzing all $x\in\mathcal{D}_{red}$, we eventually obtain two distilled datasets, $\mathcal{D}_{dir}$ and $\mathcal{D}_{int}$, for tuning.
6.3 LoRA Fine-Tuning
To fine-tune defense models with distilled datasets, we choose the publicly available Llama 2 [61] and its 7b model to demonstrate that an open-source, low-cost, yet robust defense model can be trained. Given the limited GPU resources, we utilize a parameter-efficient fine-tuning (PEFT) method known as LoRA [60] to tailor the pre-trained Llama-2-7b for dedicated jailbreak defense. LoRA is designed to adapt large pre-trained language models with minimal computational overhead. Despite the reduction in trainable parameters, LoRA maintains competitive performance, often matching or exceeding the results of full fine-tuning approaches [60].
We allocate 80% of the samples in $\mathcal{D}_{dir}$ or $\mathcal{D}_{int}$ for fine-tuning and reserve the remaining 20% as validation sets. During the training process, we continue to use our defense prompts and queries from the collected datasets as inputs; that is, we take $P_{dir}(x)$ or $P_{int}(x)$ as input and $y$ as the label to fine-tune the Llama model.
Formally, the fine-tuning objective for the direct prompt can be formulated as follows:
[…]
where $\Theta$ represents the trainable parameters of LoRA, $\Phi_{0}$ denotes the pre-trained weights of the Llama model, and $\Delta\Phi(\Theta)$ indicates the parameter increment determined by LoRA. Fine-tuning the Llama model with this objective can enhance its ability to directly detect jailbreak portions in query prompts.
Likewise, the fine-tuning objective for the intent prompt can be written as follows:
$\max_{\Theta}\sum_{(x,y)\in\mathcal{D}_{int}}\sum_{t=1}^{|y|}\log\left(p_{\Phi_{0}+\Delta\Phi(\Theta)}(y_{t}\mid P_{int}[x],y_{<t})\right)$ (4)
This objective allows the fine-tuned model to identify jailbreak intentions within query prompts.
Implementation. We use LoRA to fine tune the Llama model with $r$ of 8 and $\alpha$ of 32 [60]. The initial learning rate is $10^{-3}$. The training epochs are 1 in batch size of 8.
7 Evaluation
TABLE IV: Jailbreak ASR for various defense methods.
Target Model Defense Method Optimization-based Template-based Indirect Multilingual Normal
GCG JSAA PAIR AutoDAN DrAttack Puzzler MultiJail AlpacaEval
GPT-3.5 GPT-3.5 (baseline) 0.56 0.94 0.72 0.90 0.78 0.98 0.39 0.97
Perplexity Filter [62] 0.03 1.00 1.00 1.00 1.00 1.00 1.00 0.99
SmoothLLM [51] 0.48 0.89 0.65 0.93 0.85 0.99 1.00 0.96
Llama Guard [42] 0.41 0.61 0.43 1.00 0.97 0.93 0.95 0.99
Llama Guard 2 [63] 0.14 0.50 0.37 0.15 0.89 0.64 0.55 0.99
$P_{direct}$-tuned Shadow Stack 0.08 0.11 0.14 0.07 0.78 0.07 0.74 0.96
$P_{direct}$-tuned SelfDefend 0.06 0.11 0.07 0.07 0.62 0.07 0.30 0.94
$P_{intent}$-tuned Shadow Stack 0.08 0.05 0.16 0.09 0.18 0.20 0.57 0.99
$P_{intent}$-tuned SelfDefend 0.05 0.05 0.12 0.08 0.16 0.20 0.26 0.97
Double shadow stack 0.04 0.02 0.10 0.05 0.18 0.07 0.47 0.97
Double shadow stack + GPT-3.5 0.03 0.02 0.06 0.05 0.16 0.07 0.19 0.95
GPT-4 GPT-4 (baseline) 0.08 0.16 0.33 0.19 0.74 0.90 0.08 0.97
Perplexity Filter [62] 0.03 1.00 1.00 1.00 1.00 1.00 1.00 0.99
SmoothLLM [51] 0.07 0.19 0.33 0.18 0.91 0.88 1.00 0.97
Llama Guard [42] 0.41 0.61 0.43 1.00 0.98 0.93 0.95 0.99
Llama Guard 2 [63] 0.14 0.50 0.35 0.15 0.91 0.64 0.55 0.99
$P_{direct}$-tuned Shadow Stack 0.04 0.11 0.17 0.07 0.71 0.12 0.71 0.96
$P_{direct}$-tuned SelfDefend 0.01 0.07 0.15 0.05 0.58 0.07 0.06 0.94
$P_{intent}$-tuned Shadow Stack 0.08 0.03 0.19 0.07 0.18 0.19 0.56 0.99
$P_{intent}$-tuned SelfDefend 0.04 0.02 0.19 0.06 0.17 0.14 0.04 0.97
Double shadow stack 0.02 0.02 0.12 0.04 0.18 0.12 0.47 0.97
Double shadow stack + GPT-4 0.01 0.02 0.12 0.04 0.17 0.07 0.04 0.95
Llama 2 Llama-2-7b-chat (baseline) 0.57 0.96 0.59 0.68 0.88 0.99 0.14 0.98
Perplexity Filter [62] 0.00 1.00 1.00 1.00 1.00 1.00 1.00 0.99
SmoothLLM [51] 0.93 1.00 0.82 0.95 0.98 1.00 0.99 0.99
Llama Guard [42] 0.40 0.41 0.46 1.00 0.84 0.93 0.95 0.99
Llama Guard 2 [63] 0.17 0.53 0.41 0.15 0.89 0.64 0.55 0.99
(7b-chat) $P_{direct}$-tuned Shadow Stack 0.06 0.06 0.25 0.05 0.36 0.09 0.73 0.97
$P_{direct}$-tuned SelfDefend 0.04 0.06 0.22 0.05 0.31 0.09 0.10 0.95
$P_{intent}$-tuned Shadow Stack 0.11 0.02 0.24 0.08 0.15 0.22 0.58 0.99
$P_{intent}$-tuned SelfDefend 0.09 0.02 0.21 0.07 0.15 0.22 0.06 0.98
Double shadow stack 0.05 0.02 0.18 0.04 0.12 0.04 0.48 0.97
Double shadow stack + Llama 2 0.03 0.02 0.16 0.04 0.11 0.04 0.06 0.95
In this section, we extensively evaluate SelfDefend with Llama-2-tuned defense models and compare their performances with other jailbreak defense methods. We generally follow the setup mentioned in §5 unless explicitly specified in §7.1. For each evaluation target, we assess not only the defense effectiveness (in §7.2) and the extra delay $\Delta d$ (in §7.3) as in §5, but also whether the detected harmful portion actually aligns with the original prompt (i.e., explainability in §7.4), and how robust our defense models are against targeted optimization-based attacks (in §7.5) and prompt injection (in §7.6).
7.1 Experimental Setup
Target LLMs, Benchmarks, and Environment. As shown in Table IV, our tuned defense models are paired with protect three popular LLMs: GPT-3.5, GPT-4, and the open-source Llama-2-7b-chat [61]. We use the same benchmark datasets as in Table II (see §5). We follow the same procedure and configuration as in §5 to generate the tested jailbreak prompts for GPT-3.5 and GPT-4. The process is similar for testing the target model Llama-2-7b-chat; we simply set the target model for query-based JSAA, PAIR, and DrAttack jailbreaks to Llama-2-7b-chat to generate their jailbreak prompts. Our evaluation is implemented using PyTorch 2.2.1 and conducted on an NVIDIA TESLA H800 GPU.
Baselines. We compare our framework with popular jailbreak defense methods, including Perplexity Filter [62], SmoothLLM [51], Llama Guard [42], and Llama Guard 2 [63]. Specifically, Perplexity Filter leverages a Llama-2-7b model to calculate the perplexity of the input prompt. A jailbreak is considered to happen when the perplexity exceeds a threshold. We set this threshold at the maximum perplexity of any prompt in the JailbreakBench dataset of harmful behavior prompts. SmoothLLM perturbs the jailbreak prompts with character-level changes to enable the target LLM to perform defense. In this paper, we set SmoothLLM to perform character swapping with a 10% perturbation percentage. Llama Guard is a fine-tuned Llama-2-7b model designed to detect the toxicity category of input prompts. Llama Guard 2 is a fine-tuned version of Llama-3-8B [81].
Metric. We measure the performance of defense methods by the attack success rate (ASR), i.e., the frequency with which jailbreak prompts in a benchmark dataset bypass the detection of a defense method. The lower the ASR, the stronger the defense performance. However, the definition of “attack success” is different for different defense methods. For Perplexity Filter, the success denotes that the perplexity of the query text is no more than the fixed threshold. For SmoothLLM, the definition of success is the success of the attack method, i.e., ASR [29] or Unsafe Rate [34]. It is worth noting that the Unsafe Rate refers to the unsafety of the target model’s response, rather than the input prompt. For Llama Guard, it is success when its response is “unsafe”. For our fine-tuned models, the success represent their response being “No.”
Figure 5: An example of tuned models against GCG on GPT-3.5.
Figure 6: An example of tuned models against PAIR on GPT-3.5.
Figure 7: An example of tuned models against DrAttack on GPT-3.5.
Figure 8: An example of tuned models against MultiJail on GPT-3.5.
Figure 9: Tuned models’ response to a normal prompt from AlpacaEval.
7.2 Defense Effectiveness
We now discuss the defense effectiveness of our proposed technique. In line with Table IV, we discuss the results over GPT-3.5, GPT-4, Llama-2-7b-chat, and a “double shadow stack” setting, respectively.
Results on GPT-3.5. Table IV reports ASRs of multiple defense methods under various types of jailbreak attacks. The defense methods are evaluated across different jailbreak techniques, consisting of optimization-based, template-based, indirect, multilingual, and normal prompts. Under the optimization-based GCG and JSAA, ASR of SelfDefend with the intent prompt is 0.05, slightly better than that with the direct prompt. In contrast, SelfDefend with the direct prompt is slightly better than SelfDefend with the intent prompt under the template-based PAIR and AutoDAN.
For indirect DrAttack, ASR of the fine-tuned model with direct prompt is 0.78, the same as the original GPT-3.5, whereas the dedicated model with intent prompt reduces ASR to 0.18. With manual exploration, we find that DrAttack splits the sentence components of the original jailbreak prompt and replaces them with harmless words, making it difficult for SelfDefend (direct prompt) to extract harmful content. Although Puzzler is also an indirect attack, the prompts it produces contain offensive content as implicit clues for jailbreak, resulting in a lower ASR of SelfDefend with the direct prompt. For multilingual MultiJail, the intent prompt tuned model is superior than the direct prompt by 17%. Therefore, the model fine-tuned with the direct prompt is well-suited for identifying overt jailbreak intentions within sentences, whereas the model fine-tuned with the intent prompt excels in detecting jailbreak attempts that encompass more subtle or concealed intentions. Besides, SelfDefend has ASRs that are comparable with all other defense methods under normal prompts, indicating its compatibility with normal users.
Figures 5, 6, 7, 8 and 9 present several example responses of the tuned models toward either jailbreak prompts or a normal query. Note that in Figure 7, DrAttack evades the detection of the direct prompt tuned model, while the intent tuned model accurately identifies the intent of DrAttack, which further illustrates the weakness of the direct tuned model against indirect attacks. Moreover, as expected, we observe that both tuned models make satisfactory decisions on other jailbreaks or the normal prompt.
Results on GPT-4. Similar with the results achieved over GPT-3.5, we can observe that our defense frameworks achieve the lowest ASR under all jailbreak methods, and is still close to ASR of GPT-4 on AlpacaEval. We interpret the results as generally encouraging, illustrating the defensive superiority of SelfDefend without intruding normal user experience. Under the optimization-based GCG, the lowest ASR (as achieved by SelfDefend) is 0.01 with the direct prompt. For JSAA, SelfDefend with the intent prompt is slightly better than SelfDefend with the direct prompt. Under the template-based PAIR and AutoDAN, SelfDefend with the direct prompt is slightly superior than that with the intent prompt. For indirect DrAttack, ASR of the fine-tuned model with direct prompt is 0.71, close to that of the original GPT-4, while the fine-tuned model with intent prompt reduces ASR to 0.18. In addition, the direct prompt tuned model performs the lowest ASR under Puzzler, but far inferior to the intent one under MultiJail.
Results on Llama-2-7b-chat. Our defense frameworks exhibit the lowest ASRs across all jailbreak methods. Moreover, it achieves a ASR that is comparable to Llama-2-7b-chat on AlpacaEval, showing SelfDefend’s defensive effectiveness without compromising normal user experience. Specifically, under the white-box GCG, the ASRs for both of our SelfDefend frameworks are less than 10%.
Perplexity Filter exhibits the lowest ASR against GCG; with further exploration, we find that this is because its filter threshold is set to the maximum perplexity among all prompts in JailbreakBench, making it trivial to detect GCG with a garbled suffix. Yet, its detectability in other settings are not as good as SelfDefend. Under the optimization-based JSAA attack, SelfDefend with the intent prompt achieves the lowest ASR of 0.02, performing marginally better than with the direct prompt. In the context of template-based attacks such as PAIR and AutoDAN, the results between SelfDefend with the direct prompt and the intent one are close, with each having their own superior cases. For the indirect DrAttack, while the direct prompt tuned model largely reduces the ASR compared to the original Llama-2-7b-chat model, it still has a 21% gap compared to the intent one. Additionally, the direct prompt fine-tuned model has the lowest ASR under the Puzzler attack, yet it is notably outperformed by the intent prompt one in the MultiJail scenario.
Double Shadow Stacks. The current design of SelfDefend  accompanies one shadow stack with one LLM under protection. Despite the encouraging protection results, one may wonder incorporating multiple shadow stacks to harvest increased defense capability. Here, we define a “double shadow stack” setting as two shadow stacks independently analyze an input, where an input is deemed jailbreak when at least one shadow stack flags it as jailbreak.
In Table IV, “Double shadow stack” is instantiated by incorporating one $P_{direct}$-based and one $P_{intent}$-based shadow stacks. We can observe that the results of double shadow stack are close to the best results of the two SelfDefend frameworks, except under MultiJail. And when we further fuse the target model (i.e., normal stack) with the double shadow stack, it exhibits the lowest ASRs under all attack scenarios, as shown in Table IV. Nevertheless, we point out that when fusing two or three stacks together, the performance over normal prompts (AlpacaEval) decrease, meaning a few more normal inputs are deemed harmful. This is expected; taking the logical disjunction of multiple stacks’ predictions reduces the false negatives of SelfDefend, yet likely increases the false positives. In practice, we suggest using this paradigm only when the standard form of SelfDefend offers low detection accuracy.
Figure 10: Average $\Delta d$ introduced by the tuned models on Llama-2-7b-chat.
TABLE V: The CLIP-Score [64, 82, 83] ($\uparrow$ indicates better) based on Table IV’s results for Llama 2.
Tuned Model Text Optimization-based Template-based Indirect Multilingual
GCG JSAA PAIR AutoDAN DrAttack Puzzler MultiJail
$P_{direct}$-tuned model Generated Attack Prompts 0.85 0.70 0.83 0.78 0.74 0.66 0.68
Identified Harmful Prompts 0.95 0.94 0.90 0.94 0.79 0.73 0.69
$P_{intent}$-tuned model Generated Attack Prompts 0.85 0.70 0.83 0.78 0.73 0.67 0.67
Identified Harmful Intentions 0.88 0.88 0.85 0.88 0.83 0.77 0.74
7.3 Extra Delay $\Delta d$
Similar to how we measured the extra delay $\Delta d$ introduced by GPT-based SelfDefend in §5.3, we now report the average delay $\Delta d$ introduced by the tuned models on Llama-2-7b-chat in Figure 10. Compared to earlier results in Figure 3, $\Delta d$ under AlpacaEval is negligible for SelfDefend with both $P_{direct}$ and $P_{intent}$, with the former around 0 seconds and the latter around 0.01 seconds. In contrast, GPT-3.5-based (GPT-4-based) SelfDefend with $P_{intent}$ incurred an average $\Delta d$ of around 0.07 seconds (0.02 seconds) under AlpacaEval. This is likely because the fine-tuned models enjoy faster inference speeds when dealing with similar queries. Besides the normal prompts, the tuned models also significantly reduce the extra delay $\Delta d$ under various jailbreak prompts. For example, the maximum $\Delta d$ now decreases from the earlier 1.1 seconds on GPT-4 to 0.25 seconds. Indeed, except for JSAA, $\Delta d$ in all attack scenarios are now below 0.1 seconds, while there was no single $\Delta d$ below the same time limit for GPT-4-based SelfDefend. These results indicate that the tuning-based SelfDefend achieves negligible delays for both normal and jailbreak prompts, making it feasible for real-world deployment.
7.4 Explainability
Figure 5, 6, 7, and 8 demonstrate the efficacy of the tuned models in identifying harmful content within jailbreak prompts. Here, we further quantitatively assess whether the identified harmful content actually aligns with the original jailbreak prompt. We measure their semantic similarity using the ensemble CLIP-score [64, 82, 83] as shown in Table V. Note that our measurement exclusively targets the examples where the defense mechanism is successful.
We calculate the CLIP-score between the original prompts $P_{query}$ from JailbreakBench and the attack-generated jailbreak prompts, as well as the CLIP-score between $P_{query}$ and our identified harmful content (prompts/intentions). Specifically, the CLIP-scores for identified harmful portions are higher than those for the generated attack prompts, suggesting that the jailbreak prompts have altered the original prompts, whereas the harmful portions identified by our models are more aligned with the original prompts. Regarding the differences between the two tuned models, we observe that the similarity of identified harmful prompts is higher than that of harmful intentions in the context of optimization-based and template-based attacks, but lower than that of intentions under indirect and multilingual attacks. This can be attributed to the fact that optimization-based and template-based attack prompts contain the original prompts, whereas indirect and multilingual attack prompts significantly distort the original prompts. In summary, these findings are consistent with our initial design intentions and demonstrate that the harmful portions identified by our models can indeed provide an explainable foundation for defense mechanisms.
TABLE VI: ASR with the regular GCG attack on Llama 2 and the targeted GCG directly on our tuned defense models.
Tuned Model Regular GCG Targeted GCG
(from Table IV) (white box)
$P_{direct}$-tuned model 0.06 0.07
$P_{intent}$-tuned model 0.11 0.09
7.5 Robustness to Targeted GCG Attacks
In earlier evaluation with optimization-based attacks such as GCG in §7.2, we used a regular setting where such attacks leverage gradients to optimize adversarial suffixes on an open-source substitute model (i.e., Llama-2-7b-chat) and transfers these suffixes to attack our fine-tuned defense models. To stress-test the robustness of our fine-tuned defense models against targeted GCG attacks, we give the adversary full white-box access to our model parameters and allow them to optimize adversarial suffixes directly on our defense models.
Table VI presents the ASR results for both $P_{direct}$-tuned model and $P_{intent}$-tuned model. We observe that the ASR results of the targeted, white-box GCG remain below 10% for both types of tuned models in the shadow stack. ASR slightly increases for the targeted GCG under the $P_{direct}$-tuned model but slightly decreases for the targeted GCG under the $P_{intent}$-tuned model. This indicates that our defense models are robust to even targeted GCG attacks.
7.6 Robustness to Prompt Injection
Since SelfDefend uses $P_{direct}$ or $P_{intent}$ to wrap the original $P_{query}$, it is natural to wonder whether SelfDefend is robust to prompt injection [65, 66], another major LLM threat that is orthogonal to jailbreak attacks.
To measure the robustness of SelfDefend’s fine-tuned models against prompt injection, we leverage HouYi [66], an effective black-box prompt injection methodology that has been tested on various LLM-integrated applications. In our evaluation, we assume that HouYi [66] can directly access the outputs of our tuned defense models. HouYi’s goal is to search for a distraction prompt in an input of the form “[Jailbreak Prompt] [Distraction Prompt]” so that our defense models tend to answer with “No.” We use 100 original jailbreak goals defined in JailbreakBench as “[Jailbreak Prompt]” to conduct this testing. The ASR results for the direct and intention prompt-tuned models are 0.21 and 0.17, respectively. We consider these ASR values low for both tuned models, indicating their robustness against distractions from prompt injection. Moreover, we notice that the intent prompt-tuned model is less affected by prompt injection than the direct prompt-tuned model, as evidenced by its lower ASR.
8 Conclusion
We have introduced SelfDefend, a robust, low-cost, and self-contained defense against LLM jailbreak attacks. Inspired by the concept of shadow stacks, SelfDefend delivers a dual-layer defense mechanism comprising a shadow LLM that guards the target LLM. It further leverages a tuning-based approach to enhancing the shadow LLM’s defense capability. Evaluation shows that SelfDefend is lightweight and effective to mitigate a wide spectrum of jailbreak attacks while rarely undermining normal queries.
Appendix A Keywords
TABLE VII: The refusal signals considered for GPT-3.5/4.
“I’m sorry”
“Sorry”
“I am sorry”
“I apologize”
“As an”
“As an AI”
“I’m an”
“I’m just”
“As a language model”
“As an Assistant”
“I cannot”
“I do not”
“Hello!”
“is illegal and unethical”
“I’m really sorry”
“OpenAI” “unethical”
“not ethical”
“illegal”
“not legal”
“My apologies”
“but I cannot”
“If you have any other non-malicious requests”
“I’m not able to provide”
“I am really sorry”
“I cannot create”
“I cannot support”
“I can’t provide”
“I can’t assist”
TABLE VIII: The refusal signals considered for Llama 2.
“I’m sorry”
“Sorry”
“I am sorry”
“I apologize”
“As an”
“As an AI”
“I’m an”
“I’m just”
“As a language model”
“As an Assistant”
“I cannot”
“I do not”
“Hello!”
As mentioned in §5.2, we use a list of refusal keywords adopted by common practice [29]. They are listed in Table VII for GPT-3.5/5 and in Table VIII for Llama 2, respectively.
Appendix B More Examples from Our Tuned Models
More examples from our tuned defense models are presented in Figure 1112, and 13.
Figure 11: An example of tuned models against JSAA on GPT-3.5.
Figure 12: An example of tuned models against AutoDAN on GPT-3.5.
Figure 13: An example of tuned models against Puzzler on GPT-3.5.